# Characterizing the Sensitivity to Individual Bit Flips in Client-Side Operations of the CKKS Scheme

### v1

@@ VI Conclusions
This paper has presented a comprehensive evaluation of CKKS’s sensitivity to single‐bit flips, emphasizing the influence of RNS and NTT optimizations on error resilience. While the vanilla CKKS implementation exhibits substantial innate tolerance to individual bit‐flip faults, the optimized variants—with RNS or NTT enabled—display markedly increased vulnerability, potentially undermining the correctness of decrypted outputs.
Our findings indicate that, despite its performance drawbacks, vanilla CKKS offers a robustness advantage by minimizing the impact of isolated bit errors compared to its optimized counterparts. Furthermore, when employing large ring dimensions for security, encoding fewer slots (powers of two below the maximum) introduces redundancy that renders a subset of coefficients entirely insensitive to bit flips. This “gap”‐driven effect persists even in optimized configurations, highlighting a practical mechanism for enhancing resilience without altering core codec operations.
Overall, these results deepen our understanding of fault tolerance in homomorphic encryption schemes and provide actionable guidance for balancing throughput, security, and error resilience when selecting CKKS parameters.
Future Work
Promising directions include studying error tolerance in application‐specific domains—such as deep neural networks for image processing, where limited accuracy loss is acceptable—and conducting a detailed investigation of error propagation under multiple concurrent bit flips. Such analyses will further illuminate strategies for error mitigation in practical homomorphic computing scenarios.
@@ Acknowledgments
First, I would like to thank Pradip Bose, whose inspiration to explore fault tolerance in homomorphic encryption significantly shaped this work. I also appreciate the open‐source homomorphic encryption community and the developers of the OpenFHE and SEAL libraries, whose tools and documentation greatly facilitated our experiments. In particular, I extend my gratitude to Ahmad Al Badawi for his help setting up OpenFHE correctly and for his clear explanations and references on CKKS encoding and decoding.

### v3

@@ VI Conclusions
This paper presents a comprehensive evaluation of CKKS’s sensitivity to single‐bit flips, emphasizing the influence of RNS and NTT optimizations on error resilience. While the vanilla CKKS implementation exhibits substantial inherent tolerance to individual bit‐flip faults, the optimized variants–with RNS or NTT enabled–show markedly increased vulnerability, potentially undermining the correctness of decrypted outputs.
We provide a high‑level comparison of the different CKKS configurations evaluated in this work. For each scenario, we highlight the key resilience pattern and any notable observations.
Vanilla CKKS
Least‐significant bits exhibit robustness; error grows exponentially for higher bits up to modulus $q_{0}$. Coefficient $N/2$ robust in HEAAN 1.0 but not in OpenFHE vanilla (due to Gaussian noise injection).
Varying $\Delta$
Larger $\Delta$ shifts the onset of exponential error growth towards more significant bits. $\Delta\geq 2^{30}$ yields robust outcomes in over 80% of trials.
Effective Representation Range
Increasing integer‐part range raises catastrophic frequency and reduces robust frequency. Behavior mimics that of increasing $\Delta$.
Fault tolerance based on unused input elements
Encoding fewer elements increases the number of robust coefficients in $c_{0}$ or plaintext. Not using NTT increases substantially the computational cost.
Our findings indicate that, despite its performance drawbacks, vanilla CKKS offers a robustness advantage by minimizing the impact of isolated bit errors compared to its optimized counterparts. Furthermore, when employing large ring dimensions for security, encoding fewer elements (powers of two below the maximum) introduces redundancy that renders a subset of coefficients robust to bit-flips. This gap‐driven effect persists even in some optimized configurations, highlighting a practical mechanism for enhancing resilience at the cost of performance.
These results provide actionable guidance for balancing throughput, security, and error resilience when selecting CKKS parameters. They also lay the groundwork for future research in the context of application-specific error tolerance–e.g., deep neural networks for image processing–where limited accuracy loss is acceptable, as well as evaluation in multi-bit error scenarios.
@@ Acknowledgments
We acknowledge the support from the open‐source homomorphic encryption community and the developers of the OpenFHE and HEAAN libraries. We extend our gratitude to Dr. Ahmad Al Badawi for his help setting up OpenFHE and clear explanations. This work is partially supported by grants from Universidad de Buenos Aires (UBA) UBACYT 20620190100001BA, and CONICET PIP 11220220100027CO.